# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v2

@@ D.1.2 Categorization Examples
D.1.3 Agreement with Human Annotators
@@
D.2 Arbitrary Question Generation
Below is the prompt used for generating questions using Llama and GPT-4o.
@@ D.2 Arbitrary Question Generation
Appendix E Additional Results
E.1 Question Selection
The tables below present a breakdown of the test and validation results by the All Spans, Different Spans and Same Spans tasks, and by New Item, New Participant and New Item & Participant evaluation regimes aggregated across 10 cross-validation splits.
Model performance is compared to the Majority class baseline using a linear mixed effects model. In R notation: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. Significant gains over this baseline are marked with ’*’ $p<0.05$, ’**’ $p<0.01$ and ’***’ $p<0.001$ in superscript. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked in subscript with ’+’.
@@ E.1 Question Selection
Model New Item New Participant New Item & Participant
Chance / Majority Baseline $2.7\pm 0.6_{\text{+++}} 33$$2\pm 0.5_{\text{+++}} 33.$$3.3\pm 0.9_{\text{+++}}$
Question Similarity to RT-Weighted Passage $3.4\pm 0.4_{\text{+++}} 33$$3.3\pm 0.4_{\text{+++}}$ $34.9\pm 1.4_{\text{+++}}$
Reading Times Similarity to Question Paragraph Similarities $34.1\pm 0.6^{*}_{\text{+++}}$ $34.3\pm 0.6_{\text{+++}}$ $34.2\pm 1.2_{\text{+++}}$
Haller RNN (Haller et al., 2022) $40.4\pm 0.5^{***}_{\text{+++}}$ $43.3\pm 1.1^{***}_{\text{+++}}$ $40.8\pm 1.4^{***}_{\text{++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{49.0}\pm 0.8^{***}$ $\mathbf{49.9}\pm 0.7^{***}$ $\mathbf{47.8}\pm 1.5^{***}$
DalEye-LLaVA $3.6\pm 0.2_{\text{+++}} 33$$3.2\pm 0.3_{\text{+++}}$ $34.6\pm 1.1_{\text{+++}}$
DalEye-Llama $35.4\pm 0.4^{***}_{\text{+++}}$ $43.1\pm 1.6^{***}_{\text{+++}}$ $39.3\pm 1.0^{**}_{\text{+++}}$
@@
Model
New Item
Different Spans
New Participant
Different Spans
New Item & Participant
Different Spans
Chance / Majority Baseline $5.2\pm 0.7_{\text{+++}} 55$$5.7\pm 0.5_{\text{+++}}$ $53.5\pm 1.5_{\text{+++}}$
Question Similarity to RT-Weighted Passage $55.1\pm 3.7_{\text{+++}}$ $54.9\pm 3.7_{\text{+++}}$ $57.8\pm 3.6_{\text{+++}}$
Reading Times Similarity to Question Paragraph Similarities $54.8\pm 1.4_{\text{+++}}$ $4.8\pm 1.4_{\text{+++}} 55$$5.1\pm 1.7_{\text{+++}}$
Haller RNN (Haller et al., 2022) $4.9\pm 0.6^{***}_{\text{+++}} 66$$2\pm 0.9^{***}_{\text{+++}} 66.$$6.4\pm 1.2^{***}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{71.2}\pm 0.9^{***}$ $\mathbf{70.7}\pm 0.7^{***}$ $\mathbf{69.1}\pm 1.1^{***}$
DalEye-LLaVA $8.1\pm 1.2^{***}_{\text{+++}} 55$$5.7\pm 0.5_{\text{+++}}$ $58.7\pm 1.8^{*}_{\text{+++}}$
DalEye-Llama $57.4\pm 1.0^{*}_{\text{+++}}$ $65.7\pm 1.4^{***}_{\text{+++}}$ $59.9\pm 0.9^{**}_{\text{+++}}$
@@
Model
New Item
Same Span
New Participant
Same Span
New Item & Participant
Same Span
Chance / Majority Baseline $49.4\pm 0.5_{\text{{+++}}}$ $50.2\pm 0.6_{\text{{+++}}}$ $51.4\pm 1.8$
Question Similarity to RT-Weighted Passage $50.3\pm 0.5_{\text{{+++}}}$ $49.8\pm 0.8_{\text{{+++}}}$ $49.7\pm 2.3_{\text{{++}}}$
RT Similarity to Question-Word Similarities $51.0\pm 0.7_{\text{{+++}}}$ $51.0\pm 0.5_{\text{{+++}}}$ $53.1\pm 1.4$
Haller RNN (Haller et al., 2022) $51.7\pm 0.5^{*}_{\text{{+++}}}$ $52.6\pm 1.0^{*}_{\text{{+++}}}$ $50.3\pm 2.0_{\text{{+}}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{56.5}\pm 0.8^{***}$ $\mathbf{58.2}\pm 0.7^{***}$ $\mathbf{56.7}\pm 1.6$
DalEye-LLaVA $49.6\pm 0.4_{\text{{+++}}}$ $49.0\pm 0.4_{\text{{+++}}}$ $51.5\pm 1.6_{\text{{+}}}$
DalEye-Llama $50.4\pm 0.3_{\text{{+++}}}$ $4.2\pm 0.7^{***}_{\text{{+++}}} 55$$5.0\pm 1.5$
@@
Model All Spans
Diff Span
Same Span
@@
Reading times similarity to question-word similarities $34.1\pm 0.3_{\text{+++}}$ $54.3\pm 1.3_{\text{+++}}$ $51.1\pm 0.5_{\text{+++}}$
@@
Model New Item New Participant New Item & Participant
Chance / Majority Baseline $3.3\pm 0.5_{\text{+++}} 33$$3.0\pm 0.6_{\text{+++}}$ $31.7\pm 1.2_{\text{+++}}$
@@
RT Similarity to Question-Word Similarities $34.1\pm 0.6_{\text{+++}}$ $34.2\pm 0.6_{\text{+++}}$ $34.3\pm 1.3_{\text{+++}}$
Haller RNN (Haller et al., 2022) $42.9\pm 0.3^{***}_{\text{+++}}$ $45.7\pm 0.8^{***}_{\text{+++}}$ $42.4\pm 1.1^{***}_{\text{+++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{50.6}\pm\textbf{0.8}^{***}$ $\textbf{51.3}\pm\textbf{0.7}^{***}$ $\textbf{50.7}\pm\textbf{1.5}^{***}$
DalEye-LLaVA $34.3\pm 0.3_{\text{+++}}$ $34.8\pm 0.6^{*}_{\text{+++}}$ $38.5\pm 1.5^{**}_{\text{+++}}$
DalEye-Llama $35.8\pm 0.4^{**}_{\text{+++}}$ $40.8\pm 0.7^{***}_{\text{+++}}$ $39.1\pm 1.1^{***}_{\text{+++}}$
@@
Model
New Item
Different Spans
New Participant
Different Spans
New Item & Participant
Different Spans
Chance / Majority Baseline $5.8\pm 0.6_{\text{+++}} 55$$5.4\pm 0.6_{\text{+++}}$ $53.2\pm 0.9_{\text{+++}}$
Question Similarity to RT-Weighted Passage $54.6\pm 3.8_{\text{+++}}$ $54.4\pm 3.9_{\text{+++}}$ $57.3\pm 3.8_{\text{+++}}$
RT Similarity to Question-Word Similarities $54.3\pm 1.4_{\text{+++}}$ $54.4\pm 1.2_{\text{+++}}$ $53.8\pm 2.3_{\text{+++}}$
Haller RNN (Haller et al., 2022) $65.1\pm 0.8^{***}_{\text{+++}}$ $67.8\pm 1.0^{***}_{\text{+++}}$ $65.3\pm 0.7^{***}_{\text{+}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{71.6}\pm\textbf{0.6}^{***}$ $\textbf{71.6}\pm\textbf{0.7}^{***}$ $\textbf{70.4}\pm\textbf{1.9}^{***}$
DalEye-LLaVA $56.6\pm 1.1_{\text{+++}}$ $56.4\pm 0.8_{\text{+++}}$ $60.8\pm 1.9^{***}_{\text{+++}}$
DalEye-Llama $58.5\pm 1.1^{*}_{\text{+++}}$ $62.6\pm 0.9^{***}_{\text{+++}}$ $61.2\pm 0.9^{***}_{\text{+++}}$

### v3

@@ F.1.2 Categorization Examples
F.1.3 Agreement with Human Annotators
@@
F.2 Arbitrary and Text-only Question Generation
Below is the prompt used for generating questions using the Text-only Llama 3.1, Text-only GPT-4o-mini, Arbitrary Gemini 3, and Text-only LLaVA-OneVision.
@@ F.2 Arbitrary and Text-only Question Generation
Appendix G Additional Results
G.1 Goal Selection
G.1.1 Test and Validation Results across Subtasks and Evaluation Regimes
The tables below present a breakdown of the test and validation results by the All Spans, Different Spans and Same Spans tasks, and by New Text, New Participant and New Text & Participant evaluation regimes aggregated across 10 cross-validation splits.
Model performance is compared to the Majority class baseline using a linear mixed effects model. In R notation: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. Significant gains over this baseline are marked with ’*’ $p<0.05$, ’**’ $p<0.01$ and ’***’ $p<0.001$ in superscript. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked in subscript with ’+’ $p<0.05$, ’++’ $p<0.01$ and ’+++’ $p<0.001$.
@@ G.1.1 Test and Validation Results across Subtasks and Evaluation Regimes
Model New Participant New Text New Text & Participant
Chance / Majority Baseline $33.2\pm 0.5_{\text{+++}}$ $2.7\pm 0.6_{\text{+++}} 33$$3.3\pm 0.9_{\text{+++}}$
Question Similarity to RT-Weighted Passage $3.3\pm 0.4_{\text{+++}} 33$$3.4\pm 0.4_{\text{+++}}$ $34.9\pm 1.4_{\text{+++}}$
RT Similarity to Question-Word Similarities $34.3\pm 0.6_{\text{+++}}$ $34.1\pm 0.6^{*}_{\text{+++}}$ $34.2\pm 1.2_{\text{+++}}$
Haller RNN (Haller et al., 2022) $43.3\pm 1.1^{***}_{\text{+++}}$ $40.4\pm 0.5^{***}_{\text{+++}}$ $40.8\pm 1.4^{***}_{\text{++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{49.9}\pm 0.7^{***}$ $\mathbf{49.0}\pm 0.8^{***}$ $\mathbf{47.8}\pm 1.5^{***}$
DalEye-LLaVA $3.2\pm 0.3_{\text{+++}} 33$$3.6\pm 0.2_{\text{+++}}$ $34.6\pm 1.1_{\text{+++}}$
DalEye-Llama $43.1\pm 1.6^{***}_{\text{+++}}$ $35.4\pm 0.4^{***}_{\text{+++}}$ $39.3\pm 1.0^{**}_{\text{+++}}$
@@
Model New Participant Different Spans New Text Different Spans New Text & Participant Different Spans
Chance / Majority Baseline $5.7\pm 0.5_{\text{+++}} 55$$5.2\pm 0.7_{\text{+++}}$ $53.5\pm 1.5_{\text{+++}}$
Question Similarity to RT-Weighted Passage $4.9\pm 3.7_{\text{+++}} 55$$5.1\pm 3.7_{\text{+++}}$ $57.8\pm 3.6_{\text{+++}}$
RT Similarity to Question-Word Similarities $54.8\pm 1.4_{\text{+++}}$ $4.8\pm 1.4_{\text{+++}} 55$$5.1\pm 1.7_{\text{+++}}$
Haller RNN (Haller et al., 2022) $66.2\pm 0.9^{***}_{\text{+++}}$ $4.9\pm 0.6^{***}_{\text{+++}} 66$$6.4\pm 1.2^{***}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{70.7}\pm 0.7^{***}$ $\mathbf{71.2}\pm 0.9^{***}$ $\mathbf{69.1}\pm 1.1^{***}$
DalEye-LLaVA $55.7\pm 0.5_{\text{+++}}$ $58.1\pm 1.2^{***}_{\text{+++}}$ $58.7\pm 1.8^{*}_{\text{+++}}$
DalEye-Llama $65.7\pm 1.4^{***}_{\text{+++}}$ $57.4\pm 1.0^{*}_{\text{+++}}$ $59.9\pm 0.9^{**}_{\text{+++}}$
@@
Model New Participant Same Span New Text Same Span New Text & Participant Same Span
Chance / Majority Baseline $50.2\pm 0.6_{\text{{+++}}}$ $49.4\pm 0.5_{\text{{+++}}}$ $51.4\pm 1.8$
Question Similarity to RT-Weighted Passage $49.8\pm 0.8_{\text{{+++}}}$ $50.3\pm 0.5_{\text{{+++}}}$ $49.7\pm 2.3_{\text{{++}}}$
RT Similarity to Question-Word Similarities $51.0\pm 0.5_{\text{{+++}}}$ $51.0\pm 0.7_{\text{{+++}}}$ $53.1\pm 1.4$
Haller RNN (Haller et al., 2022) $52.6\pm 1.0^{*}_{\text{{+++}}}$ $51.7\pm 0.5^{*}_{\text{{+++}}}$ $50.3\pm 2.0_{\text{{+}}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{58.2}\pm 0.7^{***}$ $\mathbf{56.5}\pm 0.8^{***}$ $\mathbf{56.7}\pm 1.6$
DalEye-LLaVA $49.0\pm 0.4_{\text{{+++}}}$ $49.6\pm 0.4_{\text{{+++}}}$ $51.5\pm 1.6_{\text{{+}}}$
DalEye-Llama $54.2\pm 0.7^{***}_{\text{{+++}}}$ $0.4\pm 0.3_{\text{{+++}}} 55$$5.0\pm 1.5$
@@
Model All Spans Diff Span Same Span
@@
RT Similarity to Question-Word Similarities $34.1\pm 0.3_{\text{+++}}$ $54.3\pm 1.3_{\text{+++}}$ $51.1\pm 0.5_{\text{+++}}$
@@
Model New Participant New Text New Text & Participant
Chance / Majority Baseline $3.0\pm 0.6_{\text{+++}} 33$$3.3\pm 0.5_{\text{+++}}$ $31.7\pm 1.2_{\text{+++}}$
@@
RT Similarity to Question-Word Similarities $34.2\pm 0.6_{\text{+++}}$ $34.1\pm 0.6_{\text{+++}}$ $34.3\pm 1.3_{\text{+++}}$
Haller RNN (Haller et al., 2022) $45.7\pm 0.8^{***}_{\text{+++}}$ $42.9\pm 0.3^{***}_{\text{+++}}$ $42.4\pm 1.1^{***}_{\text{+++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{51.3}\pm\textbf{0.7}^{***}$ $\textbf{50.6}\pm\textbf{0.8}^{***}$ $\textbf{50.7}\pm\textbf{1.5}^{***}$
DalEye-LLaVA $34.8\pm 0.6^{*}_{\text{+++}}$ $34.3\pm 0.3_{\text{+++}}$ $38.5\pm 1.5^{**}_{\text{+++}}$
DalEye-Llama $40.8\pm 0.7^{***}_{\text{+++}}$ $35.8\pm 0.4^{**}_{\text{+++}}$ $39.1\pm 1.1^{***}_{\text{+++}}$
@@
Model New Participant Different Spans New Text Different Spans New Text & Participant Different Spans
Chance / Majority Baseline $5.4\pm 0.6_{\text{+++}} 55$$5.8\pm 0.6_{\text{+++}}$ $53.2\pm 0.9_{\text{+++}}$
Question Similarity to RT-Weighted Passage $54.4\pm 3.9_{\text{+++}}$ $54.6\pm 3.8_{\text{+++}}$ $57.3\pm 3.8_{\text{+++}}$
RT Similarity to Question-Word Similarities $54.4\pm 1.2_{\text{+++}}$ $54.3\pm 1.4_{\text{+++}}$ $53.8\pm 2.3_{\text{+++}}$
Haller RNN (Haller et al., 2022) $67.8\pm 1.0^{***}_{\text{+++}}$ $65.1\pm 0.8^{***}_{\text{+++}}$ $65.3\pm 0.7^{***}_{\text{+}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{71.6}\pm\textbf{0.7}^{***}$ $\textbf{71.6}\pm\textbf{0.6}^{***}$ $\textbf{70.4}\pm\textbf{1.9}^{***}$
DalEye-LLaVA $56.4\pm 0.8_{\text{+++}}$ $56.6\pm 1.1_{\text{+++}}$ $60.8\pm 1.9^{***}_{\text{+++}}$
DalEye-Llama $62.6\pm 0.9^{***}_{\text{+++}}$ $58.5\pm 1.1^{*}_{\text{+++}}$ $61.2\pm 0.9^{***}_{\text{+++}}$